OpenStudy (anonymous):

Given \cos(\alpha)=\frac{8}{9} and 0<\alpha<\pi/2, find the exact value of \hbox{} \cos(\alpha/2)

OpenStudy (anonymous):

\( \cos(\alpha)=\frac{8}{9}\) and \(0<\alpha<\frac{\pi}{2}\) find the exact value of \(\cos(\frac{\alpha}{2})\) just trying to make it legible

OpenStudy (anonymous):

is that the question? if so, you need the "half angle" formula do you know it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\cos(\frac{\alpha}{2})=\pm\sqrt{\frac{1+\cos(\alpha)}{2}}\] plug in \(\frac{8}{9}\) where you see \(\cos(\alpha)\) and you get it

OpenStudy (anonymous):

oh also note that since you are in quadrant 1, doubling it will be in quadrant 2, so your cosine is positive i.e. ignore the \(\pm\) and just use \(+\)

OpenStudy (anonymous):

so from there I just plug in the numberS? Thank you

OpenStudy (anonymous):

yes, you might want to simplify the compound fraction inside the radical after you write \[\cos(\frac{\alpha}{2})=\sqrt{\frac{1+\frac{8}{9}}{2}}\]

OpenStudy (anonymous):

I do not think I am getting the right answer..